OpenStudy (eiwoh):

Simplify 4 radical 6 over 5 radical 6 6 1/5 6 9/20 6 1/20 6 5/4

OpenStudy (eiwoh):

@emilee02 @KJSaif

OpenStudy (emilee02):

um

OpenStudy (emilee02):

i would say D

KjSaif (kjsaif):

could you explain @emilee02

OpenStudy (eiwoh):

PLZ LOL don't want to sound rude but i need to be certain

OpenStudy (emilee02):

OK

OpenStudy (emilee02):

I GTG I HAVE A CLASS

OpenStudy (eiwoh):

someone else then? thanks emilee

OpenStudy (misty1212):

HI!!

OpenStudy (emilee02):

WHAT GRADE ARE YOU IN HOWIE

OpenStudy (misty1212):

is it \[\frac{4\sqrt{6}}{5\sqrt{6}}\]?

OpenStudy (eiwoh):

yes it is

OpenStudy (misty1212):

ok then cancel the \(\sqrt6\) top and bottom, get \(\frac{4}{5}\)

OpenStudy (eiwoh):

yes ok. i understand that

OpenStudy (misty1212):

i think the question is something else can you post a screen shot?

OpenStudy (eiwoh):

no. sorry but its exactly how you showed it. I just need to simplify it

OpenStudy (eiwoh):

plz help this is 2nd to last question :(